What is the LCM of 16x2y5z4 and 21xy3z2?
Answer by stanbon(75887)

What is the LCM of 16x2y5z4 and 21xy3z2?
---------------
16 = 2^4
21 = 3*7
-----
LCM = (21*16)x^2y^5z^4
-----------------------------
Cheers,
Stan H.
============